OpenStudy (anonymous):

An object is dropped from a platform100 ft high. Ignoring wind resistance, what will its speed be when it reaches the ground?

OpenStudy (amistre64):

do you have a formula to play with?

OpenStudy (amistre64):

since gravity is the only thing accelerating it: a = -g but acceleration is defined as the rate of change of velocity dv/dt = -g when we integrate this to vind v: v = -gt + k, but since there is no velocity at t=0, k=0 v(t) = -gt --------------------- but how to determine when it hits the ground? we need to know its position velocity is defined as the rate of change of position ds/dt = -gt integrate again s = -1/2 gt^2 + c , when t=0 s=100 soo, c=100 s(t) = -1/2 gt^2 + 100, solve for t when s=0 0 = -1/2 gt^2 + 100

OpenStudy (mrnood):

@amistre64 It doesn't ask for t it asks for v @mathdumdum123 The formula you need is \[v ^{2}=u ^{2}+2as\] In your case v is the final velocity a is the acceleration = g = 32ft/s^2 s = distance = 100 Use the formula and those values to find v

OpenStudy (amistre64):

in my approach, i find t in order to determine v at t

OpenStudy (amistre64):

since i dont have the formulas memorized, i take a mathical approach.

OpenStudy (amistre64):

0 = -1/2 gt^2 + 100 1/2 gt^2 = 100 t^2 = 200/g t = sqrt(200/g) ------------- v = g sqrt(200/g) v = sqrt(200g)